# T2Vid: Translating Long Text into Multi-Image is the Catalyst for Video-LLMs

### v1

@@ Abstract
The success of Multimodal Large Language Models (MLLMs) in the image domain has garnered wide attention from the research community. Drawing on previous successful experiences, researchers have recently explored extending the success to the video understanding realms. Apart from training from scratch, an efficient way is to utilize the pre-trained image-LLMs, leading to two mainstream approaches, i.e. zero-shot inference and further fine-tuning with video data. In this work, our study of these approaches harvests an effective data augmentation method. We first make a deeper inspection of the zero-shot inference way and identify two limitations, i.e. limited generalization and lack of temporal understanding capabilities. Thus, we further investigate the fine-tuning approach and find a low learning efficiency when simply using all the video data samples, which can be attributed to a lack of instruction diversity. Aiming at this issue, we develop a method called T2Vid to synthesize video-like samples to enrich the instruction diversity in the training corpus. Integrating these data enables a simple and efficient training scheme, which achieves performance comparable to or even superior to using full video datasets by training with just 15% the sample size. Meanwhile, we find that the proposed scheme can boost the performance of long video understanding without training with long video samples. We hope our study will spark more thinking about using MLLMs for video understanding and curation of high-quality data. The code is released at https://github.com/xjtupanda/T2Vid.
@@ 1 Introduction
The past few years have seen the rapid progress of Multimodal Large Language Models (MLLMs) [40, 24, 1]. Apart from solving traditional vision tasks (such as VQA), these models also excel in following user instructions and generalizing to new tasks. A mainstream paradigm for developing such models takes a two-stage training strategy. The first stage, pretraining, mainly serves to align vision modality with text and inject various kinds of visual knowledge into the model. In this stage, large-scale datasets of text-image pairs are often used, such as LAION [30] and CC [32], comprising a large proportion of the total compute and injecting abundant vision knowledge into models. Some methods also incorporate OCR and detection-related data to improve foundational capabilities [6, 1]. The second stage, instruction fine-tuning, adapts models to accommodate various tasks and helps generalize to new instructions. Training in this stage typically involves instruction data obtained from self-instruction or adaptation of task-specific datasets (e.g. VQA and chart understanding datasets).
Researchers have recently shifted their focus from single-image models to more advanced ones that support video understanding. Borrowing successful experience from developing image models, some video counterparts are typically trained from scratch, following a similar two-stage training paradigm [7, 18]. Apart from this path, some researchers utilize pre-trained image-LLMs instead. Typical approaches include zero-shot inference [16, 35, 12] and further fine-tuning [22, 18, 15, 7]. Nevertheless, a further inspection of both approaches is still lacking, leaving many underlying factors that influence model performance largely unexplored. To unveil these factors, this work investigates these two approaches. Through a chain of explorations, we finally harvest an efficient scheme for fine-tuning image-LLMs for video understanding, dubbed T2Vid.
We start with the probing of zero-shot inference with Image-LLMs. Through quantitative and qualitative evaluation, we find that: (1) Due to a limited effective context window, it is challenging for image-LLMs to generalize to longer videos by simply using more frames; (2) Instead of understanding the temporal correlations among frames, image-LLMs might adopt an “attend-aggregate” pattern for video understanding. Thus, we switch to investigating the fine-tuning approach since further enhancing temporal understanding capabilities can boost the performance of video understanding. Nevertheless, we find a low learning efficiency problem through a simple data scaling experiment. An inspection of data characteristics suggests that this might be due to a lack of instruction diversity in the training corpus. To address this issue, we propose an efficient scheme to synthesize video-like samples. Specifically, we use existing text instruction data whose sample comprises a (long-context, instruction, answer) triplet. The long-context part is split into multiple segments and then further transformed into images, and the instruction and answer stay intact. Processed in this way, the synthetic samples have the same structure as video instruction data and can be incorporated seamlessly into the training corpus. Moreover, the proposed synthetic method can effectively enrich the instruction diversity in an economical way, thus boosting learning efficiency.
Comprehensive experiments demonstrate that our methods can facilitate efficient fine-tuning of image-LLMs for general video understanding and assist models in the comprehension of long videos. We hope that our findings and proposed scheme can motivate more thinking about MLLMs in video understanding tasks and spark the curation of higher-quality video data.
@@ 2.1 Multimodal Large Language Models
The research on Multi-modal large language models (MLLMs) has garnered wide attention from both industry and academia. Stimulated by the extraordinary capabilities that GPT-4 [29] series have shown, researchers have delved into developing open-sourced models that can compete with or surpass this amazing close-sourced product.
@@
Video-LLMs. Following the success of image-LLMs, subsequent endeavors aim to expand the triumph of image comprehension to more intricate video understanding. Works like Video-ChatGPT [27], VTimeLLM [13], PLLaVA [34] and LLaVA-NeXT-Video [44] attempt further fine-tune image-LLMs to enhance video understanding capability. Other research [22, 18, 15, 7] explores training from pre-trained LLM, following the basic alignment-then-finetuning paradigm similar to image-LLM. These approaches usually involve joint training that mixes image and video data in the training corpus. In this study, we build upon pre-trained image-LLMs and enhance video understanding capabilities through fine-tuning.
2.2 Zero-shot Inference for Video Understanding
Apart from further training, an alternative way is to design training-free approaches to perform zero-shot inference with image-LLMs. For example, IG-VLM [16] arranges multiple video frames into a single image grid and designs corresponding prompts to help models understand the grid format. Given a limited total resolution of image input, there is a trade-off between resolutions and the number of sampled video frames. SF-LLaVA [35] scales up the number of input frames by introducing a dual-stream design, where the slow branch utilizes more spatial tokens extracted with a lower frame rate, while the fast branch is the opposite. Free Video-LLM [12] focuses on efficient inference and introduces independent prompt-guided sampling strategies for spatial and temporal dimensions.
@@ 2.2 Zero-shot Inference for Video Understanding
However, these methods are generally evaluated on more traditional benchmarks like MSVD-QA [33], TGIF-QA [14] and ActivityNet-QA [41]). These benchmarks are generally domain-specific and focus on certain basic skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage (especially in longer videos) and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning.
@@
Recently, with the rise of benchmarks specifically designed for MLLMs [9, 18, 26], a more in-depth and comprehensive evaluation has become more accessible. Compared to previous traditional benchmarks, these newly developed benchmarks are generally more challenging, often entailing composite skills and a finer-grained understanding of the video (e.g. the plot in the movie or causal relationships between events), and can be much longer in duration (e.g. up to 60 minutes in the Video-MME benchmark). In this work, we identify the potential limitations of zero-shot inference using these newly developed video benchmarks.
3 Model Formulation
We focus on MLLMs of the mainstream token-level fusion architecture [40], which typically comprises a vision encoder, a connector, and an LLM backbone.
Given a user query and visual input, the vision encoder extracts features from the visual input, and the connector further projects the vision features to align with text embedding space. Then, the processed vision features are sent together with the text embeddings of user queries into the LLM backbone to output natural language response in an auto-regressive way:
$p(w_{o}|\textrm{w}_{V},\textrm{w}_{T})\sim\prod_{t=1}^{L}P(w_{t}|w_{<t},\textrm{w}_{V},\textrm{w}_{T})$ (1)
where $\textrm{w}_{o}=\{w_{o,t}\}_{t=1}^{L}$ is the output word token sequence of length $L$, $\textrm{w}_{V}$ represents the visual tokens processed by a vision encoder and a connector, $\textrm{w}_{T}$ is the text embeddings of the user query.
@@ 3 Model Formulation
During our exploration, we mainly utilize three image-LLMs, including Mini-InternVL-Chat-4B-V1.5 [6] (termed as InternVL hereafter), MiniCPM-Llama3-8B-V2.5 [37] (termed as MiniCPM-8B hereafter), and Idefics3-8B-Llama3 [17] (termed as Idefics3-8B hereafter). These instruction-tuned models are trained with massive image data and equipped with strong foundational capabilities. To support higher-resolution vision input, these models adopt the patchifying technique [38, 21, 23] with a dynamic resolution scheme, where an image can be cropped into multiple sub-images according to different aspect ratios. Specifically, InternVL supports up to 13 sub-images where each one is converted into 256 visual tokens; MiniCPM-8B slices images into a maximum of 10 patches, where each is represented by 96 visual tokens; Idefics3-8B supports up to 26 sub-images, each comprising 169 tokens.
4 Evaluation Setup
To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [9], MVBench [18], and TempCompass [26] as our benchmarks. We mainly use Video-MME in our probing since it comprehensively covers different visual domains and temporal ranges. We do not use traditional video-QA benchmarks (e.g. MSVD-QA [33], TGIF-QA [14], ActivityNet-QA [41]) since these benchmarks are generally domain-specific and focus on certain skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel. We illustrate more about the benchmarks used as follows:
Video-MME is a comprehensive benchmark designed for the evaluation of video-LLMs. For temporal coverage, videos of short length (up to 2 minutes), medium length (4–15 minutes), and longer duration (30–60 minutes) are included. The videos and annotations are manually collected and filtered. We only use the raw frames without the subtitles to evaluate video understanding capabilities.
TempCompass focuses on fine-grained temporal aspects, such as action, speed, and attribute change. The videos and metainformation are manually collected, after which annotations are generated by LLMs with the aid of metainformation. We use the multiple-choice QA (MCQ) format to align with other benchmarks.
To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment. More details about the implementation of this evaluation scheme are available in the Appendix.
5 Motivation
5.1 The Limitations of Zero-shot Reasoning
Methods LLM Params (B) Short Video Medium Video Long Video Overall
  Video-LLaVA [22] 7 45.3 38.0 36.2 39.9
VideoChat2 [18] 7 48.3 37.0 33.2 39.5
Chat-UniVi-v1.5 [15] 7 45.7 40.3 35.8 40.6
VideoLLaMA 2 [7] 8 56.0 45.4 42.1 47.9
InternVL (Zero-shot) [6] 3.8 61.3 51.8 44.3 52.5
VITA [10] 8x7 65.9 52.9 48.6 55.8
Kangaroo [25] 8 66.1 55.3 46.6 56.0
  InternVL (200K full data) 3.8 66.7 54.2 48.1 56.3
InternVL (30K hybrid data) 3.8 67.0 53.7 49.3 56.7
Table 1: Accuracy comparisons of different methods on the Video-MME benchmark, ranking in ascending order regarding overall performance. Our methods use only 15% of the total sample size compared to the full volume video datasets (200K) for fine-tuning and achieve comparable performance. 30K hybrid data comprise 20K data sampled from the full video datasets and 10K synthesized from text data using our T2Vid method. Bold digits indicate the best performance. Lightgray denotes zero-shot inference with image-LLM.
5.1.1 Lack of Frame Count Generalization
A look at zero-shot performance. To gain an initial grasp of zero-shot performance, we compare the results with some representative video-LLMs of similar LLM parameter size, including Video-LLaVA [22], VideoChat2 [18], Chat-UniVi-v1.5 [15], VideoLLaMA 2 [7], VITA [10], and Kangaroo [25], as summarized in Tab. 1.
The table shows that, through zero-shot inference, the image-LLM model already outperforms a variety of video-LLMs of larger LLM parameter sizes. This suggests that a well-trained image-LLM can somehow generalize to video understanding. The zero-shot video understanding capability can be attributed to the widely used training technique “patchifying”, which crops a large image into multiple sub-images and exempts the need for large-scale fine-tuning of vision encoders. Processed in this way, an image can be viewed as a series of sub-images, and the model processes all these sub-images together to understand the overall image. Similarly, a video is naturally a sequence of image frames. Due to this similarity in input structure, it is reasonable that image-LLMs somehow generalize to more frames without further fine-tuning.
Figure 1: Zero-shot performance on Video-MME when inference with a different number of frames. Image-LLMs can generalize to input more frames than seen in the training process but to a rather limited extent, making it hard to truly understand longer videos. The consistent increase of input frames brings limited gains to the performance or can even be detrimental. “max training frames” denotes the max number of patches used for a single image when training; “max training context” denotes the maximum context window set when training; “max LLM context” denotes the context window of the original LLM backbone before further multimodal training.
The influence of different frame counts. Now that image-LLMs can achieve competitive performance, we further probe the generalization potential w.r.t. input frames. Specifically, we experiment with different frame counts and visualize the changes in performance, as shown in Fig. 1. In the figure, we can observe that the performance generally first improves as more frames are available (around the green dotted line), then stays relatively stable or gradually drops (before hitting the red dotted line). The phenomenon indicates that though image-LLMs can generalize to more image numbers than seen in the training process, there is potentially an upper limit of frame number that models can effectively process. Therefore, zero-shot inference for longer video understanding can be challenging. More results and analyses in this regard are available in the Appendix.
5.1.2 Lack of Temporal Understanding Capabilities
Figure 2: A qualitative case using the prompt “Describe this video.”, where the video centers around the introduction of a trip to a town. Compared with the fine-tuned model, the zero-shot case reasons with and describes video frames in a relatively independent way without understanding the temporal correlations. This pattern follows the common behavior of the image model. The keyframes are highlighted with colors (except for red, which is used to indicate hallucinations). The numbers on the upper-left corner of each frame indicate frame order. Zoom in for better visibility.
Since image-LLMs are not trained with any video data, it is intriguing to investigate how these models can reason with videos. This prompts us to further probe if the image-LLM really understands a video and the temporal structure or just attends to multiple video frames and reasons with the information extracted from these frames, just like dealing with a patchified image. Our case study indicates that the latter case (which we call “attend-aggregate” pattern) is more likely. Specifically, we sample some videos and prompt the model in a zero-shot way, such as “Describe this video.”. A typical qualitative case, including the answers of zero-shot inference and fine-tuned model, is shown in Fig. 2.
As shown in the figure, though the model answers in a tone as if it is viewing a video, it does not really understand the temporal relationships between frames but instead aggregates information from different frames in a relatively independent way. More specifically, the text in green (corresponding to the 4th and 8th frames) mixes together the subtitle information. However, the two descriptions actually have inherent temporal structure, where the 4-th frame gives an overall introduction (of the town “Hallstatt”) and the 8-th frame introduces specific scenic spots. For the purple part (the 5th frame), the model mistakenly takes the description of transportation in town as the video’s title. Notably, though part of the upper half of the text has not shown up completely, the model successfully infers the complete text using language prior, demonstrating the model’s strong OCR capabilities. Then, in the second paragraph, the model goes back to describe the earlier 2-th and 3-th frame, respectively, using phrases like “in the background” and “another image” as if it is processing multiple images rather than a holistic video. This is expected since, when dealing with a patchified image, the model often attends to multiple sub-images, extracts information from them, and finally assembles a final answer rather than focusing on the temporal structure within these patches.
In contrast, the fine-tuned video model learns more temporal relationships and thus presents a more coherent description of frame sequence, using correct phrases as connections such as “the video starts with” and “transitions to”. Nevertheless, we notice that hallucinations [19, 39] also increase. Besides inherent noise in the training corpus, the phenomenon is highly correlated with the data bias. For example, many answers of samples in the ShareGemini dataset start with phrases like “The video begins with a black screen”, contributing to a strong prior in the fine-tuned model. We list more qualitative cases and analyses in the Appendix.
@@ 5.1.2 Lack of Temporal Understanding Capabilities
Figure 3: Video length statistics of ShareGemini and Video-ChatGPT datasets. Both datasets mostly cover videos shorter than 3.5 minutes. We extract video frames at an FPS of 1. For better visibility, we pick samples with frame numbers lower than 99.9 percentile for visualization.
5.2 The Pitfalls in Fine-tuning
In light of the limitations of zero-shot inference with image-LLMs, we turn to investigate the fine-tuning approach for improving the video understanding capabilities of models. In this section, we first illustrate the training settings and then the experimental findings.
@@ 5.2.1 Training Datasets
With 8 NVIDIA A800 GPUs, we train end-to-end (except for InternVL-4B, where we freeze the vision encoder) with a learning rate of 5e-6 and a global batch size of 64. During our investigation, we utilize two representative types of datasets, i.e., video-caption pairs and video instruction data. Specifically, we choose the ShareGemini [31] dataset and the Video-ChatGPT [27] dataset as caption and instruction data, respectively. We extract video frames at an FPS of 1 for each video. In consideration of efficiency, we use up to 64 frames for InternVL-4B and 24 frames for both MiniCPM-8B and Idefics3-8B. When the total number of frames exceeds the thresholds, we uniformly downsample the video frames. We show the statistics of video lengths in Fig. 3 and provide a more detailed introduction of the two datasets below.
@@
Video-ChatGPT. The video instruction dataset contains 100K video-instruction pairs. The videos in this collection are derived from ActivityNet [4]. The dataset’s duration is more diverse and uniform, yet the average video length is no more than 3.5 minutes. There are broadly three types of instructions: video summarization, questions about video content, and creative/generative tasks.
@@
The dataset is annotated in a semi-automatic manner. Some of the data samples are manually annotated by human annotators by refining and enriching the video captions. Other instruction data are generated by GPT-3.5 with the aid of off-the-shelf dense prediction and captioning models.
5.2.2 Low Learning Efficiency of Fine-tuning
Setting S M L Overall
Zero-shot 61.3 51.8 44.3 52.5
30K sampled data 66.2(+4.9) 53.3(+1.5) 47.4(+3.1) 55.7(+3.2)
200K full data 66.7(+0.5) 54.2(+0.9) 48.1(+0.7) 56.3(+0.6)
Table 2: Accuracy comparisons of different settings on the Video-MME benchmark, ranking in ascending order regarding overall performance. S, M, and L indicate short, medium, and long videos, respectively. Zero-shot indicates direct inference with the image-LLM, and full data setting indicates fine-tuning with all the video data samples in the training corpus. The performance gains of the latter two settings are calculated relative to the previous row.
We experiment with fine-tuning with only parts of the video data and with the full volume. Specifically, we randomly sample 15% of the total video samples from each video dataset and compare the performance as in Tab. 2. As shown in the table, compared with the short video set, the gains in medium and long video sets are relatively smaller. This can be attributed to the fact that the video lengths in the training corpus are mostly shorter than 3.5 minutes, rendering a large gap in the temporal range between training and testing (in the evaluation phase, video lengths range from 4–15 minutes and 30–60 minutes for medium and long videos, respectively).
More importantly, the additional gain of scaling up training volume from 30K to 200K data samples is limited (+0.6 overall). In view of the huge difference in data samples used, the learning efficiency of fine-tuning with these video datasets can be quite limited. The phenomenon also suggests that there could be high redundancy in the training corpus.
5.2.3 A Look at Instruction Diversity

### v3

# Sparrow: Data-Efficient Video-LLM with Text-to-Image Augmentation

@@ Abstract
Recent years have witnessed the success of Multimodal Large Language Models (MLLMs) in the vision understanding domain. The success of these models can largely be attributed to the dominant scaling law, which states that larger parameter sizes and data volumes contribute to better performance. Notably, data scaling has mainly been powered by automatic data pipelines, which center around the self-instruction of LLMs. The paradigm has been taken for granted for quite some time, but the study of the effectiveness of scaling with these data has been neglected for a long time. In this context, this work revisits scaling with synthetic data and focuses on developing video-LLMs from a data-centric perspective. Our main study approach is fine-tuning pre-trained image-LLMs with video data and investigating learning efficiency through data scaling. Results from our preliminary experiments reveal a low learning efficiency phenomenon when simply scaling up video data samples, which, through our probing, can be ascribed to a lack of instruction diversity. Aiming at this issue, we propose a data augmentation method called Sparrow, which synthesizes video-like samples from pure text instruction data. Mixing these synthetic samples with the video data enables a more efficient training scheme. Through comprehensive experiments, we demonstrate that our proposed method achieves performance comparable to or even superior to baselines trained with many more samples. Meanwhile, we find that incorporating these synthetic samples can boost the performance of long video understanding without training with long video data. The code and data examples are available at this link.
Figure 1: Performance comparison with different schemes of data scaling. We introduce Sparrow, a data-efficient training method for video-LLMs that achieves “Train less, gain more” in both general video and long video understanding performance. 0 sample indicates zero-shot inference with the image-LLM [1].
@@ 1 Introduction
The past few years have seen the rapid progress of Multimodal Large Language Models (MLLMs) [2, 3, 4]. Apart from solving traditional vision tasks (such as VQA), these models also excel in following user instructions and generalizing to new tasks. A mainstream paradigm for developing such models takes a two-stage training strategy. The first stage, pretraining, mainly serves to align vision modality with text and inject various kinds of visual knowledge into the model. In this stage, large-scale datasets of text-image pairs are often used, such as LAION [5] and CC [6], comprising a large proportion of the total compute and injecting abundant vision knowledge into models. Some methods also incorporate OCR and detection-related data to improve foundational capabilities [7, 4]. The second stage, instruction fine-tuning, adapts models to accommodate various tasks and helps generalize to new instructions. Training in this stage typically involves instruction data obtained from self-instruction or adaptation of task-specific datasets (e.g. VQA and chart understanding datasets). Recently, researchers have shifted their focus from single-image models to more advanced ones that support video understanding. Borrowing successful experience from developing image models, some video counterparts are typically trained from scratch, following a similar two-stage training paradigm [8, 9]. Apart from this path, some researchers utilize pre-trained image-LLMs instead. Typical approaches include zero-shot inference [10, 11, 12] and further fine-tuning [13, 9, 14, 8].
Notably, the success of these models can be largely ascribed to the formidable scaling law, which puts emphasis on scaling up parameter size or data volume for better model performance. For the data aspect, the scaling has mainly been driven by automatic data engines, which synthesize massive amounts of data without human labor. Nevertheless, the characteristics of learning from these synthesized video data stand out as a critical yet underexplored topic. Thus, in this work, we investigate the learning characteristics of video-LLM more deeply from a data scaling perspective. Our preliminary data scaling experiments reveal a low data efficiency problem, that is, the performance gains from utilizing multiple times more data is marginal. An inspection of data characteristics suggests this might be due to a lack of instruction diversity in the training corpus. To address this issue, we propose a data augmentation method, dubbed Sparrow [^1] Inspiration taken from the swiftness of sparrows., to enrich the diversity of instruction. The basic idea is to synthesize video-like samples from textual data and mix these synthetic data with the video samples. Specifically, we use existing text instruction data whose sample comprises a (long-context, instruction, answer) triplet. The long-context part is split into multiple segments and then further transformed into images, while the instruction and answer stay intact. Processed in this way, the synthetic samples have the same structure as video instruction data and can be incorporated seamlessly into the training corpus.
Comprehensive experiments demonstrate that our methods can facilitate data-efficient fine-tuning of image-LLMs for general video understanding and assist models in the comprehension of long videos. Specifically, using the same number of training samples, our method shows clear advantages over other data schemes. It even surpasses the baselines trained with many more samples, achieving high data efficiency (Fig. 1). The contributions of this work include:
  • We investigate the fine-tuning approach for developing video-LLMs from a data perspective, and shed light on possible factors that lead to low learning efficiency.
  • We propose a data augmentation method that improves the instruction diversity of training data and facilitates a more efficient training scheme.
  • We perform comprehensive experiments to evaluate the proposed method and examine its key properties, paving the way for future research in this line.
@@ 2.1 Multimodal Large Language Models
Video-LLMs. Following the success of image-LLMs, subsequent endeavors aim to expand the triumph to more intricate video understanding. Works like Video-ChatGPT [15], VTimeLLM [16], PLLaVA [17] and LLaVA-NeXT-Video [18] attempt further fine-tune image-LLMs to enhance video understanding capability. Other research [13, 9, 14, 8] explores training from pre-trained LLM, following the basic alignment-then-finetuning paradigm similar to image-LLM. These approaches usually involve joint training that mixes image and video data in the training corpus. In this study, we build upon pre-trained image-LLMs and enhance video understanding capabilities through fine-tuning.
2.2 Evaluation of Video Understanding
@@
Early methods [10, 15] are generally evaluated on more traditional benchmarks like MSVD-QA [19], TGIF-QA [20] and ActivityNet-QA [21]). These benchmarks are generally domain-specific and focus on certain basic skills, such as action recognition and repetition count, which lack comprehensiveness in both length coverage (especially in longer videos) and skill coverage. Moreover, the questions asked often involve shallow perception without deeper reasoning.
@@
Recently, with the rise of benchmarks specifically designed for MLLMs [22, 9, 23, 24], a more in-depth and comprehensive evaluation has become more accessible. Compared to previous traditional benchmarks, these newly developed benchmarks are generally more challenging, often entailing composite skills and a finer-grained understanding of the video (e.g. the plot in the movie or causal relationships between events), and can be much longer in duration (e.g. up to 60 minutes in the Video-MME benchmark). In this work, our study adopts these newly developed video benchmarks.
2.3 Textual Data for Video Understanding
Since MLLMs are typically built upon LLMs and thus highly compatible with textual data, some works have explored utilizing pure text data to boost the performance of video understanding. Below, we outline the key ideas of these methods and their differences from our method.
  • Textual data for context expanding. Previous works have explored utilizing textual data to expand the context window of base LLMs. Specifically, in order to facilitate long video understanding, LLaMA-VID [25] and LongVA [26] incorporate long text data in fine-tuning and continued pre-training stages, respectively, to expand the context window of LLM backbones.
    Differences: In this work, we (1) adopt textual data as a data augmentation method and (2) use them in the vision form to accommodate the training format.
  • Synthetic textual data for video understanding. This line of work investigates synthesizing textual data that simulates video QA data, aiming to transfer temporal reasoning capabilities from textual training. More specifically, TOPA [27] extracts textual captions and object-level information from video frames, while T3 [28] gathers similar information from multiple different images.
    Differences: These two works seek to boost video understanding with synthetic data, while ours aims to enrich the instruction diversity of the training corpus. Moreover, our method does not require calling advanced LLM APIs to build data; instead, our method utilizes existing datasets.
3 A Probing Study of Data Scaling
To understand the scaling characteristics of training data, our study starts with fine-tuning with different sample sizes and examines the relationship between training sample size and model performance.
In this section, we introduce the study’s training and evaluation setup and then illustrate the empirical findings.
3.1 Training Setup
3.1.1 Model Setup
@@
During our exploration, we mainly utilize two image-LLMs, including Mini-InternVL-Chat-4B-V1.5 [7] (termed as InternVL hereafter), MiniCPM-Llama3-8B-V2.5 [1] (termed as MiniCPM-8B hereafter). These instruction-tuned models are trained with massive image data and equipped with strong foundational capabilities. To support higher-resolution vision input, these models adopt the patchifying technique [29, 30, 31] with a dynamic resolution scheme, where an image can be cropped into multiple sub-images according to different aspect ratios. Specifically, InternVL supports up to 13 sub-images, each of which is converted into 256 visual tokens; MiniCPM-8B slices images into a maximum of 10 patches, where each is represented by 96 visual tokens. During training and evaluation, we switch off the patchifying option for higher efficiency.
3.1.2 Training Configurations
For fairness and ease of reproduction, we follow the official implementations. More specifically, we train the whole model end-to-end (except for InternVL-4B, where we freeze the vision encoder) with a learning rate of 5e-6.
@@ 3.1.2 Training Configurations
Figure 2: Video length statistics of ShareGemini and Video-ChatGPT datasets. Both datasets mostly cover videos shorter than 3.5 minutes. We extract video frames at an FPS of 1 for each video.. For better visibility, we pick samples with frame numbers lower than 99.9 percentile for visualization.
@@ 3.1.3 Training Datasets
During our investigation, we utilize two representative types of datasets, i.e., video-caption pairs and video instruction data. Specifically, we choose the ShareGemini [32] dataset and the Video-ChatGPT [15] dataset as caption and instruction data, respectively. For each video, frames are extracted at an FPS of 1. In consideration of efficiency, we use up to 64 frames for InternVL-4B and 24 frames for MiniCPM-8B. When the total number of frames exceeds the threshold, we uniformly downsample the video frames. The statistics of video lengths are shown in Fig. 2, and we provide more introduction to the two datasets below.
@@
Video-ChatGPT. The video instruction dataset contains 100K video-instruction pairs. The videos in this collection are derived from ActivityNet [36]. The dataset’s coverage of video duration is larger, yet the average video length is no more than 3.5 minutes. There are broadly three types of instructions: video summarization, questions about video content, and creative/generative tasks.
@@
The dataset is annotated in a semi-automatic manner. A small portion of data samples are manually annotated by human annotators by refining and enriching the video captions. Other instruction data are generated by GPT-3.5 with the aid of off-the-shelf dense prediction and captioning models.
3.2 Evaluation Setup
To evaluate the model capabilities in an efficient and comprehensive way, we use Video-MME [22], MVBench [9], and TempCompass [23] as our benchmarks. We do not use traditional video-QA benchmarks (e.g. MSVD-QA [19], TGIF-QA [20], ActivityNet-QA [21]) since these benchmarks are generally limited to a small coverage of domains, task types, and video lengths. Moreover, the questions asked often involve shallow perception without deeper reasoning since early models generally lack reasoning capacity, whereas recent LLM-based models excel. We illustrate more about the benchmarks used as follows:
Video-MME is a comprehensive benchmark designed for the evaluation of video-LLMs. For temporal coverage, videos of short length (up to 2 minutes), medium length (4–15 minutes), and longer duration (30–60 minutes) are included. The videos and annotations are manually collected and filtered. We only use the raw frames without the subtitles to focus on the evaluation of video understanding capabilities.
TempCompass focuses on fine-grained temporal aspects, such as action, speed, and attribute change. The videos and meta-information are manually collected, after which annotations are generated by LLMs with the aid of meta-information. We use the multiple-choice QA (MCQ) format to align with other benchmarks.
To ensure robust and efficient judging of model answers, we use a combination of exact matching and LLM matching for assessment. More details about the implementation of this evaluation scheme are available in Appendix A.1.
3.3 Main Findings
3.3.1 Low Learning Efficiency Issue
Figure 3: Scaling performance of fine-tuning InternVL with different data volumes and types on different benchmarks. We consider fine-tuning with ShareGemini, Video-ChatGPT, and a mix of both with a 1:$1:11$ : 1 sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference.